What is the center of the circle given by the equation below?
x2 - 16x + y2 - 18y = -129

Answer by stanbon(75887)

What is the center of the circle given by the equation below?
x2 - 16x + y2 - 18y = -129
----
Complete the square on the x and on the y terms:
x^2 -16x + 64 + y^2 -18y + 81 = -129+64+81
-----
Factor:
(x-8)^2 + (y-9)^2 = 16
----
Center: (8,9)
radius: sqrt(16) = 4
==========================
Cheers,
Stan H.
========================